# Topological Optimal Transport for Geometric Cycle Matching

### v1

@@ 4 Characterisation of geodesics
Geodesic properties have been extensively studied for the space of measure networks under Gromov-Wasserstein distance [43, 10, 33] and the space of persistence diagrams under Wasserstein distance [47, 6]. In this section, we derive similar results for the $\mathrm{TpOT}$ metric when $p=2$. These results are of theoretical interest, but we plan to explore practical implications in future work; for example, the geodesic structure and curvature properties can be used to study Fréchet means of ensembles of topological networks, extending ideas in [47, 10]. We begin by recalling some definitions from metric geometry—see [4] as a general reference.
@@ 4.1 Metric Geometry Concepts
Consider a metric space $(X,d)$. A geodesic between points $x,y\in X$ is defined as a path $\gamma:[0,1]\longrightarrow X$ with $\gamma(0)=x$, $\gamma(1)=y$ and such that, for every $0\leq s\leq t\leq 1$, we have $d(\gamma(s),\gamma(t))=(t-s)d(x,y)$. In fact, it suffices to show that $d(\gamma(s),\gamma(t))\leq(t-s)d(x,y)$ always holds, as the reverse inequality then follows for free—see, e.g., [7, Lemma 1.3].
@@
For $(X,d)$ a geodesic space, we say that it has curvature bounded below by zero [43, Section 4.2] if, for every geodesic $\gamma:[0,1]\longrightarrow X$ and every point $x\in X$, the following holds:
@@ 4.2 Geodesics between measure persistence diagrams
Set $\xi_{st}$ as the coupling in $\Pi_{\rm adm}(\nu_{s}^{\xi},\nu_{t}^{\xi})$ given by $\xi_{st}=(\phi_{s}\times\phi_{t})_{\#}\pi^{e}$, where we use the map
@@
$d^{\mathrm{MPD}}_{\mathrm{W},p}(\nu_{s}^{\xi},\nu_{t}^{\xi})^{p}\leq\int_{\overline{\Lambda}^{2}}\|\phi_{s}(\overline{\lambda})-\phi_{t}(\overline{\lambda})\|^{p}_{p}\xi_{st}(d\phi_{s}(\overline{\lambda})\times d\phi_{t}(\overline{\lambda}))=\int_{\overline{\Lambda}^{2}}\|\phi_{s}-\phi_{t}\|^{p}_{p}\;\;\xi_{st},$
@@
$\int_{\overline{\Lambda}^{2}}\|\phi_{s}-\phi_{t}\|^{p}_{p}\;\;\xi_{st}=(t-s)^{p}\int_{\overline{\Lambda}^{2}}\|d_{p}\|^{p}_{p}\;\;\xi=(t-s)^{p}d^{\mathrm{MPD}}_{\mathrm{W},p}(\nu_{0},\nu_{1})^{p}.$
@@
Consider two measure persistence diagrams $\nu_{0},\nu_{1}$. Take any path between them of the form $\nu_{t}=(\phi_{t})_{\#}\xi$, where $\xi$ is an optimal coupling between $\nu_{0},\nu_{1}$, and for every $(\lambda,\lambda^{\prime})\in\overline{\Lambda}\times\overline{\Lambda}$, the function $t\mapsto\phi_{t}(\lambda,\lambda^{\prime})$ traces out a geodesic between $\lambda$ and $\lambda^{\prime}$ in $(\overline{\Lambda},\|\cdot\|_{p})$. Then, the same argument as in Proposition 4 shows that $\nu_{t}$ is a geodesic between $\nu_{0},\nu_{1}$. In particular, this shows that $d^{\mathrm{MPD}}_{\mathrm{W},1}$ admits geodesics which are not of the form $\phi_{t}$.
@@ 4.3 Geodesics in the space of measure topological networks
where $\tilde{X}=X\times X^{\prime},\tilde{Y}=\left((Y\times Y^{\prime})\cup(Y\times\partial_{Y^{\prime}})\cup(\partial_{Y}\times Y^{\prime})\right)$ are as in Proposition 3 (we once again simplify notation by using expressions such as $Y\times\partial_{Y^{\prime}}$ rather than $Y\times\{\partial_{Y^{\prime}}\}$),
@@
$\mathrm{diag}_{\tilde{X}}:\tilde{X}$ $\rightarrow\tilde{X}\times\tilde{X}$
$\tilde{x}$ $\mapsto(\tilde{x},\tilde{x}).$
Similarly, set $\mathbf{1}_{\pi^{e}}=(\mathrm{diag}_{\tilde{Y}})_{\#}(\pi^{e})$, where $\mathrm{diag}_{\tilde{Y}}$ is defined analogously. Note that we have that $\xi_{st}\coloneqq(\iota_{s}\times\iota_{t})_{\#}\mathbf{1}_{\pi^{e}}=(\phi_{s}\times\phi_{t})_{\#}\pi^{e}$ as in the proof of Prop. 4.
By the same argument as in the proof of Prop. 4, we have that
@@
$\hskip 108.405pt=(t-s)^{p}\int_{\tilde{X}^{4}}|k(x_{1},x_{2})-k^{\prime}(x^{\prime}_{1},x^{\prime}_{2})|^{p}\>\mathrm{d}\mathbf{1}_{\pi^{v}}(\tilde{x}_{1},\tilde{x}_{1})\>\mathrm{d}\mathbf{1}_{\pi^{v}}(\tilde{x}_{2},\tilde{x}_{2})$
$\hskip 108.405pt=(t-s)^{p}\int_{\tilde{X}^{2}}|k(x_{1},x_{2})-k^{\prime}(x_{1}^{\prime},x_{2}^{\prime})|^{p}\>\mathrm{d}\pi^{v}(x_{1},x_{1}^{\prime})\>\mathrm{d}\pi^{v}(x_{2},x_{2}^{\prime}),$
@@
$\hskip 108.405pt=(t-s)^{p}\int_{\tilde{X}^{2}\times\tilde{Y}^{2}}|\omega(x,y)-\omega^{\prime}(x^{\prime},y^{\prime})|^{p}\>\mathrm{d}\mathbf{1}_{\pi^{v}}(\tilde{x},\overline{x})\>\mathrm{d}\mathbf{1}_{\pi^{e}}(\tilde{y},\tilde{y})$
$\hskip 108.405pt=(t-s)^{p}\int_{\tilde{X}\times(\overline{Y}\times\overline{Y^{\prime}})}|\omega(x,y)-\omega^{\prime}(x^{\prime},y^{\prime})|^{p}\>\mathrm{d}\pi^{v}(x,x^{\prime})\>\mathrm{d}\pi^{e}(y,y^{\prime}).$
@@
where $k_{t},\omega_{t},\pi^{v},\pi^{e},\tilde{X},\tilde{Y}$ are as in Prop. 2, and $\iota_{t}$ is such that the map
@@
The proof follows by adapting techniques from [43, Thm. 3.1] and [6, Thm. 10]. Let $[Q^{t}]$ be an arbitrary geodesic, with
@@
with $p_{i}:X^{0}\times X^{t_{1}}\times\cdots\times X^{1}\rightarrow X^{i}$ projection on the $i$th factor. Then, by suboptimality, we have
@@
$=\frac{1}{t}\|\omega^{0}-\omega^{t}\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{e}_{t})}+\frac{1}{1-t}\|\omega^{t}-\omega^{1}\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{e}_{t})}-\frac{1}{t(1-t)}\|(1-t)(\omega^{0}-\omega^{t})-t(\omega^{t}-\omega^{1})\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{e}_{t})},$ (28)
@@
$B$ $=\frac{1}{t}\|k^{0}-k^{t}\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{v}_{t})}+\frac{1}{1-t}\|k^{t}-k^{1}\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{v}_{t})}-\frac{1}{t(1-t)}\|(1-t)(k^{0}-k^{t})-t(k^{t}-k^{1})\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{v}_{t})}$ (29)
@@
$C$ $=\frac{1}{t}\|\iota^{0}-\iota^{t}\|^{2}_{L^{2}(\xi^{e}_{t})}+\frac{1}{1-t}\|\iota^{t}-\iota^{1}\|^{2}_{L^{2}(\xi^{e}_{t})}-\frac{1}{t(1-t)}\|(1-t)(\iota^{0}-\iota^{t})-t(\iota^{t}-\iota^{1})\|^{2}_{L^{2}(\xi^{e}_{t})}.$ (30)
@@
$\frac{1}{t}\left\|\omega^{0}-\omega^{t}\right\|_{L^{2}(\xi_{t}^{v}\otimes\xi_{t}^{e})}^{2}=2^{n}\cdot\frac{1}{k}\left\|\omega^{0}-\omega^{k2^{-n}}\right\|_{L^{2}(\xi_{t}^{v}\otimes\xi_{t}^{e})}^{2}$ $=2^{n}\cdot\frac{1}{k}\left\|\sum_{\ell=1}^{k}(\omega^{(\ell-1)2^{-n}}-\omega^{\ell 2^{-n}})\right\|_{L^{2}(\xi_{t}^{v}\otimes\xi_{t}^{e})}^{2}$
@@
$A$ $\leq 2^{n}\sum_{j=1}^{2^{n}}\|\omega^{(j-1)2^{-n}}-\omega^{j2^{-n}}\|_{L^{2}(\pi^{v}_{j}\otimes\pi^{e}_{j})}-\frac{1}{t(1-t)}\|(1-t)(\omega^{0}-\omega^{t})-t(\omega^{t}-\omega^{1})\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{e}_{t})},$ (33)
$B$ $\leq 2^{n}\sum_{j=1}^{2^{n}}\|k^{(j-1)2^{-n}}-k^{j2^{-n}}\|_{L^{2}(\pi^{v}_{j}\otimes pi^{v}_{j})}-\frac{1}{t(1-t)}\|(1-t)(k^{0}-k^{t})-t(k^{t}-k^{1})\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{v}_{t})},$ (34)
$C$ $\leq 2^{n}\sum_{j=1}^{2^{n}}\|\iota^{(j-1)2^{-n}}-\iota^{j2^{-n}}\|_{L^{2}(\pi^{e}_{j})}-\frac{1}{t(1-t)}\|(1-t)(\iota^{0}-\iota^{t})-t(\iota^{t}-\iota^{1})\|^{2}_{L^{2}(\xi^{e}_{t})}.$ (35)
@@
$d_{\rm TpOT}(Q^{0},Q^{1})^{2}$ $\leq A+B+C$
$\leq\left(\sum_{j=1}^{2^{n}}\left(\|\omega^{(j-1)2^{-n}}-\omega^{j2^{-n}}\|_{L^{2}(\pi^{v}_{j}\otimes pi^{e}_{j})}+\|k^{(j-1)2^{-n}}-k^{j2^{-n}}\|_{L^{2}(\pi^{v}_{j}\otimes\pi^{v}_{j})}+\|\iota^{(j-1)2^{-n}}-\iota^{j2^{-n}}\|_{L^{2}(\pi^{e}_{j})}\right)\right)$
$\qquad-\frac{1}{t(1-t)}\left(\|(1-t)(\omega^{0}-\omega^{t})-t(\omega^{t}-\omega^{1})\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{e}_{t})}+\|(1-t)(\omega^{0}-\omega^{t})-t(\omega^{t}-\omega^{1})\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{e}_{t})}\right.$
$\left.\hskip 216.81pt+\|(1-t)(\iota^{0}-\iota^{t})-t(\iota^{t}-\iota^{1})\|^{2}_{L^{2}(\xi^{e}_{t})}\right)$ (36)
$=d_{\rm TpOT}(Q^{0},Q^{1})^{2}$
$\qquad-\frac{1}{t(1-t)}\left(\|(1-t)(\omega^{0}-\omega^{t})-t(\omega^{t}-\omega^{1})\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{e}_{t})}+\|(1-t)(\omega^{0}-\omega^{t})-t(\omega^{t}-\omega^{1})\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{e}_{t})}\right.$
$\left.\hskip 216.81pt+\|(1-t)(\iota^{0}-\iota^{t})-t(\iota^{t}-\iota^{1})\|^{2}_{L^{2}(\xi^{e}_{t})}\right),$ (37)
where (37) follows by the assumption that $[Q^{t}]$ is a geodesic and by the optimality of $\pi^{e}_{t},\pi^{v}_{t}$. This inequality implies that $\pi_{v},\pi_{e}$ are optimal for $Q^{0},Q^{1}$, and that
$\|(1-t)(\omega^{0}-\omega^{t})-t(\omega^{t}-\omega^{1})\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{e}_{t})}$ $+\|(1-t)(\omega^{0}-\omega^{t})-t(\omega^{t}-\omega^{1})\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{e}_{t})}$
$\hskip 72.26999pt+\|(1-t)(\iota^{0}-\iota^{t})-t(\iota^{t}-\iota^{1})\|^{2}_{L^{2}(\xi^{e}_{t})}=0$

### v2

@@ 4 Characterisation of geodesics
Geodesic properties have been extensively studied for the space of measure networks with the Gromov-Wasserstein distance [54, 11, 43] and the space of persistence diagrams under Wasserstein distance [57, 7]. In this section, we derive similar results for the $\mathrm{TpOT}$ metric of Definition 3 when $p=2$. These results are of theoretical interest, but we plan to explore practical implications in future work; for example, the geodesic structure and curvature properties can be used to study Fréchet means of ensembles of topological networks, extending ideas in [57, 11]. We begin by recalling some definitions from metric geometry—see [4] as a general reference.
@@ 4.1 Metric Geometry Concepts
Consider a metric space $(X,d)$. A geodesic between points $x,y\in X$ is defined as a path $\gamma:[0,1]\longrightarrow X$ with $\gamma(0)=x$, $\gamma(1)=y$ and such that, for every $0\leq s\leq t\leq 1$, we have $d(\gamma(s),\gamma(t))=(t-s)d(x,y)$. In fact, it suffices to show that $d(\gamma(s),\gamma(t))\leq(t-s)d(x,y)$ always holds, as the reverse inequality then follows for free—see, e.g., [8, Lemma 1.3].
@@
For $(X,d)$ a geodesic space, we say that it has curvature bounded below by zero [54, Section 4.2] if, for every geodesic $\gamma:[0,1]\longrightarrow X$ and every point $x\in X$, the following holds:
@@ 4.2 Geodesics between measure persistence diagrams
Set $\xi_{st}$ as the coupling in $\Pi_{\rm adm}(\nu_{s}^{\xi},\nu_{t}^{\xi})$ given by $\xi_{st}=(\phi_{s}\times\phi_{t})_{\#}\xi$, where we use the map
@@
$d^{\mathrm{MPD}}_{\mathrm{W},p}(\nu_{s}^{\xi},\nu_{t}^{\xi})^{p}\leq\int_{\overline{\Lambda}^{2}}\|\phi_{s}(\overline{\lambda})-\phi_{t}(\overline{\lambda})\|^{p}_{p}\>\mathrm{d}\xi_{st}(\overline{\lambda})=\int_{\overline{\Lambda}^{2}}\|\phi_{s}-\phi_{t}\|^{p}_{p}\>\mathrm{d}\xi_{st},$
@@
$\int_{\overline{\Lambda}^{2}}\|\phi_{s}-\phi_{t}\|^{p}_{p}\;\;\mathrm{d}\xi_{st}=(t-s)^{p}\int_{\overline{\Lambda}^{2}}\|d_{p}\|^{p}_{p}\>\mathrm{d}\xi=(t-s)^{p}d^{\mathrm{MPD}}_{\mathrm{W},p}(\nu_{0},\nu_{1})^{p}.$
@@
Consider two measure persistence diagrams $\nu_{0},\nu_{1}$. Take any path between them of the form $\nu_{t}=(\phi_{t})_{\#}\xi$, where $\xi$ is an optimal coupling between $\nu_{0},\nu_{1}$, and for every $(\lambda,\lambda^{\prime})\in\overline{\Lambda}\times\overline{\Lambda}$, the function $t\mapsto\phi_{t}(\lambda,\lambda^{\prime})$ traces out a geodesic between $\lambda$ and $\lambda^{\prime}$ in $(\overline{\Lambda},\|\cdot\|_{p})$. Then, the same argument as in Proposition 5 shows that $\nu_{t}$ is a geodesic between $\nu_{0},\nu_{1}$. In particular, this shows that for $p=1$, $d^{\mathrm{MPD}}_{\mathrm{W},1}$ admits geodesics which are not of the form $\phi_{t}$.
@@ 4.3 Geodesics in the space of measure topological networks
where $\tilde{X}=X\times X^{\prime}$ and $\tilde{Y}=\left((Y\times Y^{\prime})\cup(Y\times\partial_{Y^{\prime}})\cup(\partial_{Y}\times Y^{\prime})\right)$ are as in Proposition 3 (and we once again simplify notation by using expressions such as $Y\times\partial_{Y^{\prime}}$ rather than $Y\times\{\partial_{Y^{\prime}}\}$), and
@@
$\mathrm{diag}_{\tilde{X}}:\tilde{X}\rightarrow\tilde{X}\times\tilde{X},\quad\tilde{x}\mapsto(\tilde{x},\tilde{x}).$
Similarly, set $\mathbf{1}_{\pi^{e}}=(\mathrm{diag}_{\tilde{Y}})_{\#}(\pi^{e})$, where $\mathrm{diag}_{\tilde{Y}}$ is defined analogously. Note that we have that $\xi_{st}\coloneqq(\iota_{s}\times\iota_{t})_{\#}\mathbf{1}_{\pi^{e}}=(\phi_{s}\times\phi_{t})_{\#}\pi^{e}$ as in the proof of Proposition 5.
By the same argument as in the proof of Proposition 5, we have that
@@
$\hskip 72.26999pt=(t-s)^{p}\int_{\tilde{X}^{4}}|k(x_{1},x_{2})-k^{\prime}(x^{\prime}_{1},x^{\prime}_{2})|^{p}\>\mathrm{d}\mathbf{1}_{\pi^{v}}(\tilde{x}_{1},\tilde{x}_{1})\>\mathrm{d}\mathbf{1}_{\pi^{v}}(\tilde{x}_{2},\tilde{x}_{2})$
$\hskip 72.26999pt=(t-s)^{p}\int_{\tilde{X}^{2}}|k(x_{1},x_{2})-k^{\prime}(x_{1}^{\prime},x_{2}^{\prime})|^{p}\>\mathrm{d}\pi^{v}(x_{1},x_{1}^{\prime})\>\mathrm{d}\pi^{v}(x_{2},x_{2}^{\prime}),$
@@
$\hskip 72.26999pt=(t-s)^{p}\int_{\tilde{X}^{2}\times\tilde{Y}^{2}}|\omega(x,y)-\omega^{\prime}(x^{\prime},y^{\prime})|^{p}\>\mathrm{d}\mathbf{1}_{\pi^{v}}(\tilde{x},\overline{x})\>\mathrm{d}\mathbf{1}_{\pi^{e}}(\tilde{y},\tilde{y})$
$\hskip 72.26999pt=(t-s)^{p}\int_{\tilde{X}\times(\overline{Y}\times\overline{Y^{\prime}})}|\omega(x,y)-\omega^{\prime}(x^{\prime},y^{\prime})|^{p}\>\mathrm{d}\pi^{v}(x,x^{\prime})\>\mathrm{d}\pi^{e}(y,y^{\prime}).$
@@
where $k_{t},\omega_{t},\pi^{v},\pi^{e},\tilde{X},\tilde{Y}$ are as in Proposition 6, and $\iota_{t}$ is such that the map
@@
The proof follows by adapting techniques from [54, Theorem 3.1] and [7, Theorem 10]. Let $[Q^{t}]$ be an arbitrary geodesic, with
@@
with $p_{i}:X^{0}\times X^{t_{1}}\times\cdots\times X^{1}\rightarrow X^{i}$ projection on the $i$th factor. Then, by sub-optimality, we have
@@
$=\frac{1}{t}\|\omega^{0}-\omega^{t}\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{e}_{t})}+\frac{1}{1-t}\|\omega^{t}-\omega^{1}\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{e}_{t})}$
$\hskip 108.405pt-\frac{1}{t(1-t)}\|(1-t)(\omega^{0}-\omega^{t})-t(\omega^{t}-\omega^{1})\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{e}_{t})},$ (31)
@@
$\begin{split}B&=\frac{1}{t}\|k^{0}-k^{t}\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{v}_{t})}+\frac{1}{1-t}\|k^{t}-k^{1}\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{v}_{t})}\\ &\hskip 108.405pt-\frac{1}{t(1-t)}\|(1-t)(k^{0}-k^{t})-t(k^{t}-k^{1})\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{v}_{t})}\end{split}$ (32)
@@
$\begin{split}C&=\frac{1}{t}\|\iota^{0}-\iota^{t}\|^{2}_{L^{2}(\xi^{e}_{t})}+\frac{1}{1-t}\|\iota^{t}-\iota^{1}\|^{2}_{L^{2}(\xi^{e}_{t})}\\ &\hskip 108.405pt-\frac{1}{t(1-t)}\|(1-t)(\iota^{0}-\iota^{t})-t(\iota^{t}-\iota^{1})\|^{2}_{L^{2}(\xi^{e}_{t})}.\end{split}$ (33)
@@
$\frac{1}{t}\left\|\omega^{0}-\omega^{t}\right\|_{L^{2}(\xi_{t}^{v}\otimes\xi_{t}^{e})}^{2}$ $=2^{n}\cdot\frac{1}{k}\left\|\omega^{0}-\omega^{k2^{-n}}\right\|_{L^{2}(\xi_{t}^{v}\otimes\xi_{t}^{e})}^{2}$ (34)
$l$ $=2^{n}\cdot\frac{1}{k}\left\|\sum_{\ell=1}^{k}(\omega^{(\ell-1)2^{-n}}-\omega^{\ell 2^{-n}})\right\|_{L^{2}(\xi_{t}^{v}\otimes\xi_{t}^{e})}^{2}$
@@
$\begin{split}A&\leq 2^{n}\sum_{j=1}^{2^{n}}\|\omega^{(j-1)2^{-n}}-\omega^{j2^{-n}}\|_{L^{2}(\pi^{v}_{j}\otimes\pi^{e}_{j})}\\ &\hskip 108.405pt-\frac{1}{t(1-t)}\|(1-t)(\omega^{0}-\omega^{t})-t(\omega^{t}-\omega^{1})\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{e}_{t})},\end{split}$ (37)
$\begin{split}B&\leq 2^{n}\sum_{j=1}^{2^{n}}\|k^{(j-1)2^{-n}}-k^{j2^{-n}}\|_{L^{2}(\pi^{v}_{j}\otimes\pi^{v}_{j})}\\ &\hskip 108.405pt-\frac{1}{t(1-t)}\|(1-t)(k^{0}-k^{t})-t(k^{t}-k^{1})\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{v}_{t})},\end{split}$ (38)
$\begin{split}C&\leq 2^{n}\sum_{j=1}^{2^{n}}\|\iota^{(j-1)2^{-n}}-\iota^{j2^{-n}}\|_{L^{2}(\pi^{e}_{j})}\\ &\hskip 108.405pt-\frac{1}{t(1-t)}\|(1-t)(\iota^{0}-\iota^{t})-t(\iota^{t}-\iota^{1})\|^{2}_{L^{2}(\xi^{e}_{t})}.\end{split}$ (39)
@@
$d_{\rm TpOT}(Q^{0},Q^{1})^{2}$
$\leq A+B+C$
$\begin{split}&\leq\sum_{j=1}^{2^{n}}\left(\|\omega^{(j-1)2^{-n}}-\omega^{j2^{-n}}\|_{L^{2}(\pi^{v}_{j}\otimes\pi^{e}_{j})}+\|k^{(j-1)2^{-n}}-k^{j2^{-n}}\|_{L^{2}(\pi^{v}_{j}\otimes\pi^{v}_{j})}\right.+\\ &\hskip 252.94499pt\|\left.\iota^{(j-1)2^{-n}}-\iota^{j2^{-n}}\|_{L^{2}(\pi^{e}_{j})}\right)\\ &\hskip 18.06749pt-\frac{1}{t(1-t)}\left(\|(1-t)(\omega^{0}-\omega^{t})-t(\omega^{t}-\omega^{1})\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{e}_{t})}\right.\\ &\hskip 90.3375pt\left.+\|(1-t)(\omega^{0}-\omega^{t})-t(\omega^{t}-\omega^{1})\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{e}_{t})}\right.\\ &\left.\hskip 162.6075pt+\|(1-t)(\iota^{0}-\iota^{t})-t(\iota^{t}-\iota^{1})\|^{2}_{L^{2}(\xi^{e}_{t})}\right)\end{split}$
$\begin{split}&=d_{\rm TpOT}(Q^{0},Q^{1})^{2}\\ &\hskip 18.06749pt-\frac{1}{t(1-t)}\left(\|(1-t)(\omega^{0}-\omega^{t})-t(\omega^{t}-\omega^{1})\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{e}_{t})}\right.\\ &\hskip 90.3375pt\left.+\|(1-t)(\omega^{0}-\omega^{t})-t(\omega^{t}-\omega^{1})\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{e}_{t})}\right.\\ &\left.\hskip 162.6075pt+\|(1-t)(\iota^{0}-\iota^{t})-t(\iota^{t}-\iota^{1})\|^{2}_{L^{2}(\xi^{e}_{t})}\right),\end{split}$ (40)
where (LABEL:eq:uniquefinal) follows by the assumption that $[Q^{t}]$ is a geodesic and by the optimality of $\pi^{e}_{t},\pi^{v}_{t}$. This inequality implies that $\pi_{v},\pi_{e}$ are optimal for $Q^{0},Q^{1}$, and that
$\|(1-t)(\omega^{0}-\omega^{t})-t(\omega^{t}-\omega^{1})\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{e}_{t})}$
$\hskip 72.26999pt+\|(1-t)(\omega^{0}-\omega^{t})-t(\omega^{t}-\omega^{1})\|^{2}_{L^{2}(\xi^{v}_{t}\otimes\xi^{e}_{t})}$
$\hskip 144.54pt+\|(1-t)(\iota^{0}-\iota^{t})-t(\iota^{t}-\iota^{1})\|^{2}_{L^{2}(\xi^{e}_{t})}=0$